\begin{thebibliography}{10}

\bibitem{Alessandrini:1990}
Giovanni Alessandrini.
\newblock Singular solutions of elliptic equations and the determination of
  conductivity by boundary measurements.
\newblock {\em J. Differential Equations}, 84(2):252--272, 1990.

\bibitem{Alessandrini:Gaburro:2009}
Giovanni Alessandrini and Romina Gaburro.
\newblock The local {C}alder\`on problem and the determination at the boundary
  of the conductivity.
\newblock {\em Comm. Partial Differential Equations}, 34(7-9):918--936, 2009.

\bibitem{Ammari:Bonnetier:Capdebocsg:Fink:Tanter:2008}
Habib Ammari, Eric Bonnetier, Yves Capdeboscq, Mickael Tanter, and Mathias
  Fink.
\newblock {Electrical Impedance Tomography by Elastic Deformation}.
\newblock {\em SIAM Journal on Applied Mathematics}, 68(6):1557--1573, June
  2008.

\bibitem{Aronsson:1996}
Gunnar Aronsson.
\newblock On $p$-harmonic functions, convex duality and an asymptotic formula
  for injection mould filling.
\newblock {\em European Journal of Applied Mathematics}, 7:417--437, October
  1996.

\bibitem{Aronsson:Evans:Wu:1996}
Gunnar Aronsson, Lawrence.C. Evans, and Y.~Wu.
\newblock Fast/slow diffusion and growing sandpiles.
\newblock {\em Journal of Differential Equations}, 131(2):304 -- 335, 1996.

\bibitem{Aronsson:Janfalk:1992}
Gunnar Aronsson and Ulf Janfalk.
\newblock On {H}ele-{S}haw flow of power-law fluids.
\newblock {\em European Journal of Applied Mathematics}, 3:343--366, December
  1992.

\bibitem{Bal:2012}
Guillaume Bal.
\newblock {Cauchy problem for Ultrasound Modulated EIT}.
\newblock January 2012.
\newblock Preprint at arXiv.

\bibitem{Bal:Schotland:2010}
Guillaume Bal and John~C Schotland.
\newblock Inverse scattering and acousto-optic imaging.
\newblock {\em Physical review letters}, 104(4), 2010.

\bibitem{Borcea:2002}
Liliana Borcea.
\newblock Electrical impedance tomography.
\newblock {\em Inverse Problems}, 18(6):R99, 2002.
\newblock With an addendum~\cite{Borcea:2003}.

\bibitem{Borcea:2003}
Liliana Borcea.
\newblock Addendum to '{Electrical impedance tomography}'.
\newblock {\em Inverse Problems}, 19(4):997, 2003.
\newblock Addendum to~\cite{Borcea:2002}.

\bibitem{Brander:Kar:Salo:2014}
Tommi Brander, Manas Kar, and Mikko Salo.
\newblock Enclosure method for the $p$-{L}aplace equation.
\newblock {\em ArXiv e-prints}, October 2014.
\newblock Available online
  \href{http://arxiv.org/abs/1410.4048}{ar{X}iv:1410.4048}.

\bibitem{Barber:Brown:Seagar:1985}
B~H Brown, D~C Barber, and A~D Seagar.
\newblock Applied potential tomography: possible clinical applications.
\newblock {\em Clinical Physics and Physiological Measurement}, 6(2):109, 1985.

\bibitem{Brown:2001}
Russell~M. Brown.
\newblock Recovering the conductivity at the boundary from the {D}irichlet to
  {N}eumann map: a pointwise result.
\newblock {\em J. Inverse Ill-Posed Probl.}, 9(6):567--574, 2001.

\bibitem{Calderon:1980}
Alberto~P. Calder{\'o}n.
\newblock On an inverse boundary value problem.
\newblock In W.H. Meyer and M.A. Raupp, editors, {\em Seminar on numerical
  analysis and its applications to continuum physics}, pages 65--73. Sociedade
  {Brasileira} de Matematica, 1980.
\newblock Reprinted as \cite{Calderon:2006}.

\bibitem{Calderon:2006}
Alberto~P. Calder{\'o}n.
\newblock On an inverse boundary problem.
\newblock {\em Computation and applied mathematics}, 25(2--3):133--138, 2006.
\newblock Reprint, originally published as \cite{Calderon:1980}.

\bibitem{Calderon:1961}
Alberto~Pedro Calder{\'o}n.
\newblock Lebesgue spaces of differentiable functions and distributions.
\newblock In {\em Partial differential equations}, volume~4 of {\em Proceedings
  of Symposia in Pure Mathematics}, pages 33--49, Providence, Rhode Island,
  USA, 1961. American mathematical society.

\bibitem{D'Onofrio:Iwaniec:2005}
Luigi D'Onofrio and Tadeusz Iwaniec.
\newblock Notes on {$p$}-harmonic analysis.
\newblock In {\em The {$p$}-harmonic equation and recent advances in analysis},
  volume 370 of {\em Contemporary Mathematics}, pages 25--50. American
  Mathematical Society, Providence, RI, 2005.

\bibitem{Garcia:Zhang:2012}
Andoni Garc\'ia and Guo Zhang.
\newblock Reconstruction from boundary measurements for less regular
  conductivities.
\newblock {\em arXiv}, December 2012.
\newblock With an appendix by R. Brown and the authors.

\bibitem{Gebauer:Scherzer:2008}
Bastian Gebauer and Otmar Scherzer.
\newblock Impedance-acoustic tomography.
\newblock {\em SIAM Journal on Applied Mathematics}, 69(2):565--576, 2008.

\bibitem{Gilbarg:Trudinger:1983}
David Gilbarg and Neil~S. Trudinger.
\newblock {\em Elliptic partial differential equations of second order}, volume
  224 of {\em Grundlehren der Mathematischen Wissenschaften [Fundamental
  Principles of Mathematical Sciences]}.
\newblock Springer-Verlag, Berlin, second edition, 1983.

\bibitem{Haberman:Tataru:2013}
Boaz Haberman and Daniel Tataru.
\newblock Uniqueness in {C}alder\'on's problem with {L}ipschitz conductivities.
\newblock {\em Duke Math. J.}, 162(3):496--516, 2013.

\bibitem{Hajlasz:Koskela:Tuominen:2008}
Piotr Haj{\l}asz, Pekka Koskela, and Heli Tuominen.
\newblock Measure density and extendability of {S}obolev functions.
\newblock {\em Revista Matem{\'a}tica Iberoamericana}, 24(2):645--669, April
  2008.

\bibitem{Hauer:2014}
Daniel Hauer.
\newblock The $p$-{D}irichlet-to-{N}eumann operator with applications to
  elliptic and parabolic problems.
\newblock March 2014.
\newblock Preprint, retrieved 18.9.2014 from
  \url{www.maths.usyd.edu.au/u/pubs/publist/preprints/2014/hauer-9.html}.

\bibitem{Heinonen:Kilpelainen:Martio:1993}
Juha Heinonen, Tero Kilpel{\"a}inen, and Olli Martio.
\newblock {\em Nonlinear potential theory of degenerate elliptic equations}.
\newblock Oxford Mathematical Monographs. The Clarendon Press, Oxford
  University Press, New York, Oxford, 1993.
\newblock Oxford Science Publications.

\bibitem{Hoell:Moradifan:Nachman:2013}
Nicholas Hoell, Amir Moradifam, and Adrian Nachman.
\newblock Current density impedance imaging of an anisotropic conductivity in a
  known conformal class.
\newblock 2013.
\newblock Preprint at arXiv:1302.7044.

\bibitem{Henkelman:Joy:Scott:1989}
M.~Joy, G.~Scott, and M.~Henkelman.
\newblock In vivo detection of applied electric currents by magnetic resonance
  imaging.
\newblock {\em Magnetic Resonance Imaging}, 7(1):89 -- 94, 1989.

\bibitem{Kang:Yun:2002}
Hyeonbae Kang and Kihyun Yun.
\newblock Boundary determination of conductivities and {R}iemannian metrics via
  local {D}irichlet-to-{N}eumann operator.
\newblock {\em SIAM J. Math. Anal.}, 34(3):719--735 (electronic), 2002.

\bibitem{Karhunen:2013}
Kimmo Karhunen.
\newblock {\em Electrical resistance tomography imaging of concrete}.
\newblock PhD thesis, {University of Eastern Finland}, 2013.

\bibitem{Kim:Kwon:Seo:Yoon:2002}
S.~Kim, O.~Kwon, J.~Seo, and J.~Yoon.
\newblock On a nonlinear partial differential equation arising in magnetic
  resonance electrical impedance tomography.
\newblock {\em SIAM Journal on Mathematical Analysis}, 34(3):511--526, 2002.

\bibitem{Kohn:Vogelius:1984}
Robert Kohn and Michael Vogelius.
\newblock Determining conductivity by boundary measurements.
\newblock {\em Communications on pure and applied mathematics}, 37(3):289--298,
  May 1984.

\bibitem{Kuijper:2007}
Arjan Kuijper.
\newblock {p-Laplacian driven image processing}.
\newblock In {\em Image Processing, 2007. ICIP 2007. IEEE International
  Conference on}, volume~5, pages V--257. IEEE, 2007.

\bibitem{Ladyzhenskaya:Ural'tseva:1968}
Olga~A. Ladyzhenskaya and N.~N. Ural'tseva.
\newblock {\em Linear and Quasilinear Elliptic Equations}.
\newblock Mathematics in Science and Engineering : a series of monographs and
  textbooks. Academic Press, 1968.

\bibitem{Lieberman:1988}
Gary~M. Lieberman.
\newblock Boundary regularity for solutions of degenerate elliptic equations.
\newblock {\em Non-Linear Analysis}, 12(11):1203--1219, November 1988.

\bibitem{Liimatainen:Salo:2012}
Tony Liimatainen and Mikko Salo.
\newblock {$n$}-harmonic coordinates and the regularity of conformal mappings.
\newblock September 2012.
\newblock Preprint at arXiv:1209.1285.

\bibitem{Lindqvist:2006}
Peter Lindqvist.
\newblock {\em Notes on the {$p$-L}aplace equation}, volume 102 of {\em Reports
  of {U}niversity of {J}yv{\"a}skyl{\"a} Department of Mathematics and
  Statistics}.
\newblock University of {J}yv{\"a}skyl{\"a}, Jyv{\"a}skyl{\"a}, Finland, 2006.

\bibitem{Iwaniec:Manfredi:1989}
Juan~Jos{\'e} Manfredi and Tadeusz Iwaniec.
\newblock Regularity of p-harmonic functions on the plane.
\newblock {\em Revista Matematica Iberoamericana}, 5(1):1--19, 1989.

\bibitem{Nachman:Tamasan:Timonov:2009}
Adrian Nachman, Alexandru Tamasan, and Alexandre Timonov.
\newblock Recovering the conductivity from a single measurement of interior
  data.
\newblock {\em Inverse Problems}, 25(3):035014, 2009.

\bibitem{Nachman:1996}
Adrian~I. Nachman.
\newblock Global uniqueness for a two-dimensional inverse boundary value
  problem.
\newblock {\em Ann. of Math. (2)}, 143(1):71--96, 1996.

\bibitem{Nakamura:Siltanen:Tamura:Wang:2005}
G.~Nakamura, S.~Siltanen, K.~Tanuma, and S.~Wang.
\newblock {Numerical Recovery of Conductivity at the Boundary from the
  Localized Dirichlet to Neumann Map}.
\newblock {\em Computing}, 75(2-3):197--213, 2005.

\bibitem{Nakamura:Tanuma:2003}
Gen Nakamura and Kazumi Tanuma.
\newblock Formulas for reconstructing conductivity and its normal derivative at
  the boundary from the localized {D}irichlet to {N}eumann map.
\newblock In {\em Recent development in theories \& numerics}, pages 192--201.
  World Sci. Publ., River Edge, NJ, 2003.

\bibitem{Radon:1913}
Johann Radon.
\newblock {Theorie und Anwendungen der absolut additiven Mengenfunktionen}.
\newblock {\em {S.-B. Akad. Wiss. Wien}}, 122:1295--1438, 1913.

\bibitem{Rellich:1940}
F.~Rellich.
\newblock {Darstellung der Eigenwerte von $\Delta u + \lambda u = 0$ durch ein
  Randintegral}.
\newblock {\em Mathematische Zeitschrift}, 46:635--636, 1940.

\bibitem{Riesz:1928a}
Fr{\'e}d{\'e}ric Riesz.
\newblock {Sur la convergence en moyenne, I}.
\newblock {\em {Acta Sci. Math (Szeged)}}, 4:58--64, 1928-1929.

\bibitem{Riesz:1928b}
Fr{\'e}d{\'e}ric Riesz.
\newblock {Sur la convergence en moyenne, II}.
\newblock {\em {Acta Sci. Math (Szeged)}}, 4:182--185, 1928-1929.

\bibitem{Salo:Zhong:2012}
Mikko Salo and Xiao Zhong.
\newblock An inverse problem for the $p$-{L}aplacian: {B}oundary determination.
\newblock {\em {SIAM} journal on mathematical analysis}, 44(4):2474--2495,
  March 2012.

\bibitem{Armstrong:Henkelman:Joy:Scott:1991}
G.C. Scott, M.~L~G Joy, R.L. Armstrong, and R.~M. Henkelman.
\newblock Measurement of nonuniform current density by magnetic resonance.
\newblock {\em Medical Imaging, IEEE Transactions on}, 10(3):362--374, 1991.

\bibitem{Siltanen:Tamminen:2011}
S.~{Siltanen} and J.~P. {Tamminen}.
\newblock {Reconstructing conductivities with boundary corrected D-bar method}.
\newblock {\em Journal of inverse and ill-posed problems}.
\newblock To be published. Preprint at arXiv:1109.5982.

\bibitem{Stein:1970}
Elias~M Stein.
\newblock {\em Singular integrals and differentiability properties of
  functions}, volume~30 of {\em Princeton mathematical series}.
\newblock Princeton university press, Princeton, New Jersey, USA, 1970.

\bibitem{Sylvester:Uhlmann:1988}
John Sylvester and Gunther Uhlmann.
\newblock Inverse boundary value problems at the boundary---continuous
  dependence.
\newblock {\em Comm. Pure Appl. Math.}, 41(2):197--219, 1988.

\bibitem{Uhlmann:2009}
Gunther Uhlmann.
\newblock Electrical impedance tomography and {C}alder{\'o}n's problem.
\newblock {\em Inverse problems}, 25(12):123011, 2009.

\bibitem{Wolff:2007}
Thomas~H. Wolff.
\newblock Gap series constructions for the {$p$-L}aplacian.
\newblock {\em Journal d'Analyse Mathematique}, 102(1):371--394, August 2007.
\newblock Preprint written in 1984.

\end{thebibliography}

\documentclass[10pt,a4paper]{amsart}
\pdfoutput=1
\usepackage[latin1]{inputenc}
\usepackage{amsmath,amssymb}
\usepackage{mathtools}
\usepackage{esint}
\usepackage{color}
\usepackage{hyperref}

\DeclareMathOperator{\dive}{div}

\DeclareMathOperator{\supp}{supp}

\DeclareMathOperator{\sign}{sign}

\newcommand*{\dx}[1]{\, d #1}
\newcommand*{\R}{\mathbb{R}}

\newcommand*{\eps}{\varepsilon}
\newcommand*{\doo}{\partial}
\newcommand*{\ol}[1]{\overline{#1}}
\newcommand*{\gradeps}{\left( |\nabla u_\eps|^2 + \eps \right)}
\newcommand{\abs}[1]{\left| #1 \right|}
\newcommand{\aabs}[1]{\left\| #1 \right\|}

\newtheorem{theorem}{Theorem}
\newtheorem{defin}[theorem]{Definition}
\newtheorem{lemma}[theorem]{Lemma}

\newtheorem{remark}[theorem]{Remark}

\numberwithin{equation}{section}
\numberwithin{theorem}{section}

\title[Calder\'on problem for the $p$-Laplacian]{Calder\'on problem for the $p$-Laplacian: First order derivative of conductivity on the boundary}
\author{Tommi Brander}
\address{Department of Mathematics and Statistics, P.O.Box 35 (MaD) FI-40014 University of Jyv\"askyl\"a, Finland}
\email{tommi.o.brander@jyu.fi}
\date{\today}

\begin{document}

\begin{abstract}
We recover the gradient of a scalar conductivity defined on a smooth bounded open set in $\mathbb{R}^d$ from the Dirichlet to Neumann map arising from the $p$-Laplace equation. For any boundary point we recover the gradient using Dirichlet data supported on an arbitrarily small neighbourhood of the boundary point. We use a Rellich-type identity in the proof. Our results are new when $p \neq 2$. In the $p = 2$ case boundary determination plays a role in several methods for recovering the conductivity in the interior.
\end{abstract}

\subjclass[2000]{
Primary
35R30, 
35J92
}

\keywords{Calder\'on's problem, inverse problems, p-laplace equation, boundary determination}

\maketitle 


\section{Introduction}
Throughout the article we assume $1 < p < \infty$, unless explicitly written otherwise.
We investigate a generalisation of the Calder\'on problem to the case of the $p$-Laplace equation:
Given a bounded open set $\Omega \subset \R^d$ and a bounded conductivity $\gamma > 0$ on $\ol\Omega$ we have the Dirichlet problem
\begin{equation} \label{eq:p-laplace}
\begin{cases}
\Delta_p^\gamma(u) = \dive \left( \gamma(x) |\nabla u|^{p-2} \nabla u \right) = 0 &\text{in $\Omega$,} \\
u = v &\text{on $\doo \Omega$}.
\end{cases}
\end{equation}
For $v \in W^{1,p}(\Omega)$ there exists a unique weak solution that satisfies $\| u \|_{W^{1,p}} \leq C \| v \|_{W^{1,p}}$, where the constant $C$ does not depend on $v$.
This is completely standard, see e.g. \cite[Proposition A.1]{Salo:Zhong:2012}.
The equation arises from the variational problem of minimising the energy $\int_\Omega \gamma |\nabla u|^p \dx{x}$.

We next define the Dirichlet to Neumann map (hereafter DN map).
We use the notation $X'$ for the dual space of continuous linear functionals on a Banach space $X$.
In the $p=2$ case the DN map gives the stationary measurements of electrical current for given voltage $v$.
\begin{defin}[Dirichlet to Neumann map]\label{definition:DN_map}
Suppose  $\Omega \subset \R^d$ is bounded open $C^1$-set and $0 < \gamma_0 < \gamma \in L^\infty(\overline{\Omega})$ for some constant~$\gamma_0$.
The weak DN map
\begin{equation*}
\Lambda_\gamma^w \colon W^{1,p}(\Omega)/W^{1,p}_0(\Omega) \to \left(W^{1,p}(\Omega)/W^{1,p}_0(\Omega)\right)'
\end{equation*}
is defined by
\begin{equation}\label{eq:weak_DN}
\left\langle \Lambda_\gamma^w(v), g \right\rangle = \int_\Omega \gamma |\nabla u|^{p-2} \nabla u \cdot \nabla \tilde g \dx{x},
\end{equation}
where $u$ solves the boundary value problem~\eqref{eq:p-laplace} with boundary values~$v$, and $\tilde g \in W^{1,p}(\Omega)$ with trace~$g$ on the boundary.
The strong DN map is defined pointwise on $\doo \Omega$ by the expression
\begin{equation}\label{eq:strong_DN}
\Lambda_\gamma^s (v)(x_0) = \gamma(x_0) |\nabla u(x_0)|^{p-2}\nabla u(x_0) \cdot \nu(x_0)
\end{equation}
when it is well-defined.
\end{defin}
We also write $\Lambda_\gamma (v)$ when the function $v$ is continuous and defined on a superset of $\doo \Omega$.
With sufficient regularity we can recover the values of the strong DN map from the weak DN map (see lemma~\ref{lemma:weak_DN_to_strong}).

Given knowledge of the DN map we determine $\nabla \gamma|_{\doo \Omega}$, which in the $p=2$ case stands for the gradient of conductivity on the boundary.
\begin{theorem}\label{thm:main}
Suppose $\Omega \subset \R^d$, $d > 1$, is a bounded open $C^{2,\beta}$ set for some $0 < \beta <1$ and the conductivities $\gamma_1$ and $\gamma_2$ are bounded from below by a positive constant and continuously differentiable with H{\"o}lder-continuous derivatives in~$\ol{\Omega}$.
If $\Lambda_{\gamma_1}^w = \Lambda_{\gamma_2}^w$, then $\nabla \gamma_1|_{\doo \Omega} = \nabla \gamma_2|_{\doo \Omega}$.
\end{theorem}
We use an explicit sequence of boundary values to reconstruct $\nabla \gamma (x_0)$ at a boundary point $x_0$.
The sequence is supported in arbitrarily small neighbourhood of $x_0$.
The boundary values we use were first introduced by Wolff~\cite{Wolff:2007} and used by Salo and Zhong~\cite{Salo:Zhong:2012} to recover the conductivity on the boundary.

Salo and Zhong~\cite{Salo:Zhong:2012} show that for all boundary points $x_0 \in \doo \Omega$ there is a sequence of solutions $(u_M)_M$ such that
\begin{equation}
M^{n-1}N^{1-p}\int_\Omega g(x) |\nabla u_M|^p \dx{x} \to c_p g(x_0)
\end{equation}
as $M, N(M) \to \infty$, with explicitly computable constant $c_p$ (see equation~\eqref{eq:c_p}).
They recover the integrals with $g = \gamma$ from the DN map.
We recover the integrals with $g = \alpha \cdot \nabla \gamma$ from a Rellich-type identity, theorem~\ref{thm:rellich}, for arbitrary direction~$\alpha \in \R^d$.

The Calder\'on problem was first introduced in~\cite{Calderon:1980}. For a review see~\cite{Uhlmann:2009}.

There have been several boundary determination results in the $p = 2$ case. Boundary uniqueness for derivatives of the conductivity was first proven by Kohn and Vogelius~\cite{Kohn:Vogelius:1984}.
Sylvester and Uhlmann in ~\cite{Sylvester:Uhlmann:1988} recovered, for smooth conductivity in smooth domain, conductivity and all its derivatives on the boundary by considering $\Lambda_\gamma$ as a pseudodifferential operator.
Nachman~\cite{Nachman:1996} recovered $\gamma|_{\doo \Omega}$ in Lipschitz domain when $\gamma \in W^{1,q}$, $q > d$, and its first derivative when $\gamma \in W^{2,q}$, $q > d/2$.
Alessandrini~\cite{Alessandrini:1990} used singular solutions with singularity near boundary to recover all derivatives of $\gamma|_{\doo \Omega}$ with less regularity assumptions on $\gamma$.
There have been several local boundary determination results, e.g.~\cite{Nakamura:Tanuma:2003,Kang:Yun:2002}.
Other reasonably recent boundary determination results include~\cite{Alessandrini:Gaburro:2009,Brown:2001}.

A Rellich identity was used by Brown, Garcia and Zhang~\cite[appendix]{Garcia:Zhang:2012} to recover the gradient of conductivity on the boundary in the $p=2$ case.
The Rellich identity was, to the best of our knowledge, introduced in~\cite{Rellich:1940}.

Several results (e.g.~\cite{Alessandrini:1990}) rely on investigating the difference $\Lambda_{\gamma_1} - \Lambda_{\gamma_2}$ of DN~maps at different conductivities.
This is difficult in the present setting due to non-linearity of the $p$-Laplace equation.
We use Rellich identity, theorem~\ref{thm:rellich}, to avoid this problem.

Electrical impedance tomography has applications in medical (e.g.~\cite{Barber:Brown:Seagar:1985}) and industrial imaging (e.g.~\cite{Karhunen:2013}), and geophysics and environmental sciences; see e.g.\ the review~\cite{Borcea:2002} and references therein.
There are practical numerical algorithms for boundary determination in the $p=2$ case, e.g.~\cite{Nakamura:Siltanen:Tamura:Wang:2005}.
Boundary determination is used in recovering the conductivity in the interior, e.g.~\cite{Haberman:Tataru:2013}.
For example, the algorithm in~\cite{Siltanen:Tamminen:2011} uses the values of conductivity and its first derivative on the boundary to extend the conductivity, thence applying to conductivities that are not constant on the boundary.

For more on the $p$-Laplace equation see e.g.~\cite{Lindqvist:2006,Heinonen:Kilpelainen:Martio:1993,D'Onofrio:Iwaniec:2005}.
The equation has applications in e.g.\ image processing~\cite{Kuijper:2007}, fluid mechanics~\cite{Aronsson:Janfalk:1992}, plastic moulding~\cite{Aronsson:1996}, and modelling of sand--piles~\cite{Aronsson:Evans:Wu:1996}.

The Calder\'on problem for the $p$-Laplace equation was first introduced in~\cite{Salo:Zhong:2012}.
The authors consider the real and the complex case separately.
In the complex case they define $p$-harmonic versions of complex geometrical optics solutions to create highly oscillating functions focused around a given boundary point and use them as Dirichlet data, thus recovering conductivity at the boundary point in question.
In the real case they replace the $p$-harmonic CGO solutions with real-valued functions having similar behaviour.
The real-valued functions were originally introduced by Wolff~\cite{Wolff:2007}.

Further progress on the $p$-Calder\'on problem was made by the author, Kar and Salo~\cite{Brander:Kar:Salo:2014}.
In the article we show that we can detect the convex hulls of inclusions, which are regions of significantly higher or lower conductivity, from the DN map.
Our main tools are the $p$-harmonic functions of Wolff and a monotonicity inequality.

Hauer~\cite{Hauer:2014} has investigated the DN map related to the $p$-Laplace equation.

One method of investigating the Calder\'on-type inverse problems for non-linear equations is based on studying the G\^ateaux derivatives of the map $\Lambda_\gamma$ at constant boundary values $a$.
In our case this does not work~\cite[appendix]{Salo:Zhong:2012}:
\begin{equation}
\Lambda_\gamma (a+tf) = t^{p-1}\Lambda_\gamma(f)
\end{equation}
for positive $t$.
For $p < 2$ the G\^ateaux derivates do not exist and for $p > 2$ the higher derivates fail to exist, or vanish, though one of them might equal $\Lambda_\gamma (f)$.
Hence nonlinear methods are necessary.

Other inverse problems related to nonlinear equations similar to the $p$-Laplace equation have been investigated before; the $1$-Laplace equation is used in current density imaging, e.g.~\cite{Kim:Kwon:Seo:Yoon:2002,Nachman:Tamasan:Timonov:2009,Hoell:Moradifan:Nachman:2013,Henkelman:Joy:Scott:1989,Armstrong:Henkelman:Joy:Scott:1991}, and the $0$-Laplace equation is related to ultrasound modulated electrical impedance tomography, e.g.~\cite{Bal:2012,Ammari:Bonnetier:Capdebocsg:Fink:Tanter:2008,Bal:Schotland:2010,Gebauer:Scherzer:2008}.

In section~\ref{sec:prelim} we introduce our notation and state lemmata.
In section~\ref{sec:salo_zhong} we restate boundary determination results we use in this article.
In the final section~\ref{section:proof} we reconstruct $\nabla \gamma|_{\doo \Omega}$ from the DN~map.

\section{Preliminaries}\label{sec:prelim}
We use the following notation:
We write $\doo_\alpha = \alpha \cdot \nabla$ for vectors $\alpha$.
On the $C^1$ boundary $\doo \Omega$ of an open set $\Omega$ we decompose the gradient of a function $f$ defined in $\ol{\Omega}$ as $\nabla f(x_0) = \nabla_\nu f(x_0) + \nabla_T f(x_0)$, where $\nabla_\nu$ is the normal component of the derivative, i.e.\ the orthogonal projection of the gradient to the normal space of the boundary, and $\nabla_T$ is the tangential derivative, i.e.\ the orthogonal projection of the gradient on the tangent space of the boundary.
We have the following identity: $|\nabla_\nu f|^2 = (\doo_\nu f)^2$.
We denote the outer unit normal by $\nu$.

We write the H{\"o}lder seminorm of order $k$ as $|f|_{k,\beta}$ and the characteristic function of a set $A$ as
\begin{equation*}
\chi_A(x) =
\begin{cases}
1 \text{ when } x \in A \\
0 \text{ when } x \notin A.
\end{cases}
\end{equation*}

We need the following elementary inequality:
\begin{lemma}\label{lemma:A.3}
Suppose $0 \leq A,B \in \R$ and $2 \leq q < \infty$.
Then for some $C > 0$ we have
\begin{equation}
|A^{q-1} - B^{q-1}| \leq C \left( A + B \right)^{q-2} | A - B|
\end{equation}
\end{lemma}
\begin{proof}
By fundamental theorem of calculus
\begin{equation*}
|A^{q-1} - B^{q-1}| = (q-1)\abs{A-B}\abs{\int_0^1 \left(A + t(B-A)\right)^{q-2} \dx{t}},
\end{equation*}
from which the estimate follows by observing that $ 0 \leq A+t(B-A) \leq A+B$.
\end{proof}

It is well-known that the solutions of the $p$-Laplace equation are in $C^{1,\beta}$. For proof of the following lemma see e.g.~\cite{Lieberman:1988}.
\begin{lemma}[Regularity result] \label{lemma:regularity}
Suppose $\Omega$ is a bounded open $C^{1,\beta_1}$ set with $0 < \beta_1 \leq 1$,
and suppose the conductivity $0 < \gamma \in C^{0,\beta_2}(\ol{\Omega})$ is bounded from above and away from zero.
Consider the weighted $p$-Laplace equation~\eqref{eq:p-laplace} with boundary values $v \in C^{1,\beta_1}(\doo \Omega)$.
Then the solution~$u$ of the weighted $p$-Laplace equation~\eqref{eq:p-laplace} is in $C^{1,\beta_3}(\ol{\Omega})$ for some $\beta_3 > 0$.
\end{lemma}
In general the solutions are not twice continuously differentiable~\cite{Iwaniec:Manfredi:1989}.

To establish Rellich identity, theorem~\ref{thm:rellich}, we use the following interpolation lemma similar to~\cite[lemma A.1]{Liimatainen:Salo:2012}:
\begin{lemma}[Interpolation lemma] \label{lemma:interpolation}
Suppose $\Omega$ is a bounded open set in $\R^d$ satisfying the measure density condition; that is, suppose there exist $\delta_0, C > 0$ such that for all $0 < \delta < \delta_0$ and for all $y \in \ol\Omega$ we have
\begin{equation}\label{eq:mdc}
|B(y,\delta)| \leq C |B(y,\delta) \cap \ol\Omega|.
\end{equation}
Also suppose $f \in C^\beta(\ol{\Omega})$ for some $\beta > 0$.
Take $1 \leq p < \infty$.
Let $M > 0$ be a constant such that $|f|_{0,\beta} \leq M$.
Then
\begin{equation}
\| f \|_{L^\infty(\ol{\Omega})} \leq C_{d,p,\beta,\Omega}M^{\frac{d}{d+\beta p}} \|f\|_{L^p(\Omega)}^\frac{\beta p}{d+ \beta p}.
\end{equation}
\end{lemma}
The measure density condition~\eqref{eq:mdc} is satisfied in a wide variety of situations~\cite{Hajlasz:Koskela:Tuominen:2008}.
In particular, domains that admit a Sobolev extension theorem satisfy the condition~\cite[proposition 1]{Hajlasz:Koskela:Tuominen:2008}.
Domains of class $C^{1,\alpha}$, and more generally Lipschitz domains, certainly admit a Sobolev extension theorem~\cite[theorem 12]{Calderon:1961}, \cite[chapter VI, section 3.1, theorem 5]{Stein:1970}.
A bounded Lipschitz set has a finite number of connected components, and thus also satisfies the condition.

The proof of the interpolation lemma is almost the same as in~\cite{{Liimatainen:Salo:2012}}.
\begin{proof}
Let $y \in \ol{\Omega}$ and write $B_\delta = B(y,\delta)$.
For $\delta > 0$ we have
\begin{align*}
\|f\|_{L^p(\ol{\Omega}\cap B_\delta)} &\geq \|f(y)\|_{L^p(\ol{\Omega}\cap B_\delta)} - \|f(x) - f(y)\|_{L^p(\ol{\Omega}\cap B_\delta)} \\
&= |f(y)|\left|\ol{\Omega}\cap B_\delta\right|^{1/p} - \left( \int_{\ol{\Omega}\cap B_\delta} |f(x) - f(y)|^p \dx{x} \right)^{1/p} \\
&\geq c_{d,p,\Omega} \delta^{d/p}|f(y)| - |f|_{0,\beta}\left(\int_{\ol{\Omega}\cap B_\delta} |x-y|^{\beta p} \dx{x}\right)^{1/p} \\
&\geq c_{d,p,\Omega} \delta^{d/p}|f(y)| - C_{d,p,\beta} \delta^{\beta + d/p}M,
\end{align*}
whereby it follows that
\begin{equation}
\| f \|_{L^\infty(\ol{\Omega})} \leq C_{d,p,\Omega}\delta^{-d/p} \|f\|_{L^p(\ol{\Omega})} + C_{d,p,\beta}\delta^\beta M.
\end{equation}
Choose
\begin{equation}
\delta = \left( \frac{\|f\|_{L^p(\ol\Omega)}}{M} \right)^\frac{p}{d + \beta p}.
\end{equation}
\end{proof}

In our proofs we use the $\eps$-perturbed $p$-Laplace equation.
Let $\eps > 0$ and define $u_\eps$ as the solution of the boundary value problem
\begin{equation} \label{eq:eps-laplace}
\begin{cases}
\dive \left( \gamma(x) \left(|\nabla u_\eps|^2 + \eps\right)^\frac{p-2}{2} \nabla u_\eps \right) = 0 &\text{in $\Omega$,} \\
u_\eps = v &\text{on $\doo \Omega$}.
\end{cases}
\end{equation}
By calculus of variations there exists a unique solution to~\eqref{eq:eps-laplace} when $\Omega$ is a bounded open set and $v \in W^{1,p}(\Omega)$.
The energy corresponding to~\eqref{eq:eps-laplace} is
\begin{equation}\label{eq:eps_energy}
\int_\Omega \gamma \left(|\nabla u_\eps|^2 + \eps\right)^{p/2} \dx{x}.
\end{equation}

\begin{lemma}[Regularity result for perturbed equation]\label{lemma:eps_regularity}
Suppose $\Omega$ is a bounded open set with $C^{1,\beta_1}$ boundary.
Suppose that boundary values $v \in C^{1,\beta_1}(\doo \Omega)$ and conductivity $\gamma$ is H{\"o}lder-continuous.
Then $u_\eps \in C^{1,\beta_2}(\ol \Omega)$ with some $0 < \beta_2 < 1$ independent of $\eps$ and the norm $|u_\eps|_{1,\beta_2}$ has an upper bound independent of~$\eps$.

Furthermore, suppose that $\Omega$ has $C^{2,\beta_3}$ boundary, $v \in C^{2,\beta_3}(\doo \Omega)$ and that $\nabla \gamma$ is H{\"o}lder-continous.
Then $ u_\eps \in C^{2,\beta_4}(\ol{\Omega})$ for some $0 < \beta_4 < 1$, which depends on~$\eps$.
\end{lemma}
\begin{proof}
That $\nabla u_\eps$ are H{\"o}lder-continuous with $\beta_2$ and the corresponding norm independent of $\eps$ follows from~\cite{Lieberman:1988}.
The $C^2$--regularity is also standard; see for example~\cite[part 4, section 8]{Ladyzhenskaya:Ural'tseva:1968} or~\cite[section 15.5]{Gilbarg:Trudinger:1983}.
\end{proof}

\begin{lemma}[Convergence of the perturbed equations]\label{lemma:perturbed_convergence}
Suppose $\Omega$ is a bounded open set with $C^{1,\beta}$ boundary, $0 < \beta < 1$, and boundary values $v$ are in $C^{1,\beta}$.
Also assume that the conductivity $\gamma$ is H\"older continuous.
Then $u_\eps \to u$ in $C^{1}(\ol \Omega)$.
\end{lemma}
\begin{proof}
We first show that the energy~\eqref{eq:eps_energy} of the perturbed equation~\eqref{eq:eps-laplace} converges to the energy of the non-perturbed equation~\eqref{eq:p-laplace}:
The solution $u$ minimises energy, and so we have
\begin{align*}
&\int_\Omega \gamma |\nabla u|^{p}\dx{x}\\
\leq &\int_\Omega \gamma |\nabla u_\eps|^{p}\dx{x}\\
\leq &\int_\Omega \gamma \left(|\nabla u_\eps|^2+\eps\right)^{p/2}\dx{x} \\
\leq &\int_\Omega \gamma \left(|\nabla u|^2+\eps\right)^{p/2}\dx{x} \\
\leq &\int_\Omega \gamma |\nabla u|^{p}\dx{x} + C
\begin{cases}
\int_\Omega \eps \left( 2|\nabla u|^2 + \eps \right)^{-1+p/2}\dx{x} &\text{ when } p > 2 \\
\int_\Omega \eps^{p/2} \dx{x} &\text{ when } p \leq 2,
\end{cases}
\end{align*}
where the latter integrals vanish as $\eps \to 0$ since $|\nabla u|$ is bounded (lemma~\ref{lemma:regularity}).
We used the fact that $u_\eps$ minimises the perturbed energy~\eqref{eq:eps_energy}, and also lemma~\ref{lemma:A.3} with $q = 1+p/2 > 2$ when $p>2$, and the fact that for non-negative numbers $(a+b)^q \leq a^q + b^q$ when $0 \leq q = p/2 \leq 1$, in the final inequality.

We established that the energy of $u_\eps$ converges to the energy of $u$ as $\eps \to 0$.
Since $\aabs{u_\eps}_{L^\infty}$ are uniformly bounded, it follows that $u_\eps$ are bounded in $W^{1,p}$ with weight.
This implies that the sequence has a weakly converging subsequence with limit that has lesser or equal energy than $u$.
But $u$ uniquely minimises energy, so $u_\eps \to u$ weakly in weighted $W^{1,p}$.
Since the weighted $L^p$ norms of $\nabla u_\eps$ converge to weighted $L^p$ norm of $\nabla u$, and the sequence converges weakly, we get that $\nabla u_\eps \to \nabla u$ strongly in weighted $L^{p}$ space; this is the Radon-Riesz property proven in~\cite{Radon:1913,Riesz:1928a,Riesz:1928b}.
Since the weighted and standard norms are equivalent, $\nabla u_\eps \to \nabla u$ strongly in standard $L^p$.

Since $\nabla u_\eps$ are uniformly H{\"o}lder in $\ol{\Omega}$ (lemma~\ref{lemma:eps_regularity}) we can use the interpolation lemma~\ref{lemma:interpolation} to deduce convergence of $\nabla u_\eps$ to $\nabla u$ in $C(\ol{\Omega})$.
The same argument holds for $u_\eps$, since by Friedrichs-Poincar\'e inequality $u_\eps \to u$ strongly in $L^p$.
\end{proof}

\section{Previous results on boundary determination for $p$-Laplacian}\label{sec:salo_zhong}
We use several results in the paper~\cite{Salo:Zhong:2012} and restate them here for convenience.

Suppose $\Omega \subset \R^d$ is an open bounded $C^1$ set and suppose $\rho \in C^1(\R^d)$ is its boundary defining function;
that is,
\begin{equation*}\Omega = \{ x \in \R^d ; \rho (x) > 0 \},\quad \doo \Omega = \{ x \in \R^d ; \rho (x) = 0 \}
\end{equation*}
and $\nabla \rho \neq 0$ on $\doo \Omega$.
By translation we may assume that we are trying to recover $\nabla \gamma$ at the origin $0 \in \doo \Omega$, and by rotation and scaling we may assume that
\begin{equation*}
-\nabla \rho (0) = \nu(0) = -e_d,
\end{equation*}
the $d$th coordinate vector.
We define the map $f \colon \Omega \to \R^d_+$ by
\begin{equation*}
f(x',x_d) = (x',\rho(x))
\end{equation*}
for $x = (x',x_d)$.
Since $\rho$ is a $C^1$ function, the map is close to the identity near the origin, and hence invertible in some neighbourhood of the origin.

We now introduce the $p$-harmonic oscillating functions of Wolff~\cite{Wolff:2007}.
\begin{lemma} \label{lemma:wolff}
Define the function $h \colon \R^d \to \R$ by $h(x)=e^{-x_d}a(x_1)$,
where $a \colon \R \to \R$ is the solution to the differential equation
\begin{equation}
a''(x_1) + V(a,a')a = 0
\end{equation}
with
\begin{equation}
V(a,a') = \frac{(2p-3)(a')^2 + (p-1)a^2}{(p-1)(a')^2 + a^2}.
\end{equation}
Then $\Delta_p^1(h)=0$ and the function $a$ is smooth and periodic with period $\lambda(p)$, so that $\int_0^\lambda a(x_1)\dx{x_1} = 0$.
\end{lemma}
For proof of the lemma see~\cite[lemma 3.1]{Salo:Zhong:2012}.

We define a smooth positive cutoff function $\zeta \in C^\infty_0(\R^d)$, such that $\zeta (x) = 1$ when $|x| < 1/2$ and $\supp \zeta \subset B(0,1)$.
We write
\begin{equation}\label{eq:v_M}
v_M(x) = h(Nf(x))\zeta(Mx),
\end{equation}
where $M$ and $N = N(M)$ are large positive numbers, $M = o(N)$.
We define $u_M$ to be the solutions of the initial value problem
\begin{equation}
\begin{cases}
\Delta_p^\gamma(u_M) = 0 &\text{in $\Omega$,} \\
u_M = v_M &\text{on $\doo \Omega$}.
\end{cases}
\end{equation}

In \cite[lemma 3.3]{Salo:Zhong:2012} Salo and Zhong show that
\begin{equation}
M^{d-1}N^{1-p}\int_\Omega \gamma |\nabla v_M|^p \dx{x} \to c_p \gamma(0)
\end{equation}
as $M \to \infty$.
The constant $c_p$ is explicit:
\begin{equation}\label{eq:c_p}
c_p = \frac{K}{p}\int_{\R^{d-1}}(\eta(x',0))^p\dx{x'}; \text{ with } K = \lambda^{-1} \int_0^\lambda \left(a^2(t) + (a'(t))^2\right)^{p/2}\dx{t}.
\end{equation}
The proof also holds when $\gamma$ is replaced by any other function continuous at~$0$, so as $M \to \infty$ we have
\begin{equation}
M^{d-1}N^{1-p}\int_\Omega \doo_\alpha \gamma |\nabla v_M|^p \dx{x} \to c_p \doo_\alpha \gamma(0).
\end{equation}
Using~\cite[lemma 3.4]{Salo:Zhong:2012} we get the following result:
\begin{lemma}\label{lemma:sz_main}
Suppose $\Omega \subset \R^d$ is a bounded open set with $C^1$ boundary, $x_0 \in \doo \Omega$, $d \geq 2$, and $g$ is continuous.
Then
\begin{equation}
M^{d-1}N^{1-p}\int_\Omega g(x) |\nabla u_M|^p \dx{x} \to c_p g(x_0)
\end{equation}
as $M \to \infty$.
\end{lemma}

\section{Proof of the main result} \label{section:proof}
In this section we establish that the quantity
\begin{equation}
\int_\Omega \doo_\alpha\gamma(x) |\nabla u|^p \dx{x}
\end{equation}
can be calculated from the measurements represented by the DN map, with the function~$u$ solving the weighted $p$-Laplace equation~\eqref{eq:p-laplace} with sufficiently smooth boundary values.
Then it follows from lemma~\ref{lemma:sz_main} that we can recover $\doo_\alpha \gamma(x_0)$ at any boundary point $x_0 \in \doo \Omega$.
The main result, theorem~\ref{thm:main}, immediately follows.

\begin{theorem}[Rellich identity] \label{thm:rellich}
Suppose $1<p<\infty$ and that $\Omega \subset \R^d$ is a bounded open set with $C^{2,\beta}$ boundary for some $0 < \beta < 1$.
Let $0 < \gamma \in C^1(\overline{\Omega})$ and let $u$ be a weak solution of the equation $\Delta_p^\gamma u = 0$.
Let $\alpha \in \R^d$.

Then
\begin{equation}
\int_\Omega (\doo_\alpha \gamma) |\nabla u|^p \dx{x} = \int_{\doo \Omega} \gamma (\alpha \cdot \nu) |\nabla u|^p \dx{S} - p\int_{\doo \Omega} (\doo_\alpha u) \gamma |\nabla u|^{p-2} \doo_\nu u \dx{S}.
\end{equation}
\end{theorem}
\begin{proof}
Suppose we have the following identity for the solution~$u_\eps$ of the perturbed $p$-Laplace equation~\eqref{eq:eps-laplace}:
\begin{align*}
&\int_\Omega (\doo_\alpha \gamma) \gradeps^{p/2} \dx{x} \\
&= \int_{\doo \Omega} \gamma (\alpha \cdot \nu) \gradeps^{p/2} \dx{S} - p\int_{\doo \Omega} \doo_\alpha u_\eps \gamma \gradeps^\frac{p-2}{2} \doo_\nu u_\eps \dx{S}.
\end{align*}
Then, when $\eps \to 0$, $\nabla u_\eps \to \nabla u$ uniformly (by lemma~\ref{lemma:perturbed_convergence}), so we get the claimed identity.

We now prove the perturbed identity, integrating by parts twice:
\begin{align*}
\int_{\doo \Omega} &(\alpha \cdot \nu) \gamma \gradeps^{p/2} \dx{S} - \int_\Omega (\doo_\alpha \gamma) \left(|\nabla u_\eps|^2 +\eps\right)^{p/2} \dx{x} \\
= & \int_\Omega \gamma \doo_\alpha \left( \gradeps^{p/2} \right) \dx{x} \\
= & p\int_\Omega \gamma \gradeps^\frac{p-2}{2} \nabla u_\eps \cdot \nabla \doo_\alpha u_\eps \dx{x} \\
= & p \int_{\doo \Omega} \doo_\alpha u_\eps \gamma \gradeps^\frac{p-2}{2} \doo_\nu u_\eps \dx{S} \\
&- p\int_\Omega \dive \left( \gamma \gradeps^\frac{p-2}{2} \nabla u_\eps \right) \doo_\alpha u_\eps \dx{x} \\
= &p\int_{\doo \Omega} \doo_\alpha u_\eps \gamma \gradeps^\frac{p-2}{2} \doo_\nu u_\eps \dx{S}.\qedhere
\end{align*}
\end{proof}

We need to recover $\gamma|_{\doo \Omega}$ and $\nabla u|_{\doo \Omega}$ to recover $\int_\Omega \doo_\alpha\gamma |\nabla u|^p \dx{x}$ from the DN~map.

Supposing $\Omega$ is an open set with $C^1$ boundary and $\gamma$ is continuous, we can recover $\gamma(x_0)$ for all boundary points $x_0$.
This follows directly from \cite{Salo:Zhong:2012}.

To calculate $\doo_\nu u$ we use $C^{1,\beta}$ boundary values $v$ so that the solution~$u$ of the boundary value problem~\eqref{eq:p-laplace} is continuously differentiable up to the boundary (by lemma~\ref{lemma:regularity}), so $\doo_\nu u$ is defined pointwise.
Recovering $\doo_\nu u$ from the strong definition of the DN map is straightforward, and we do so in lemma~\ref{lemma:doo_nu}.
However, we do not a~priori have access to the strong DN map, and must recover it from the weak definition~\ref{definition:DN_map}.
This we do in lemma~\ref{lemma:weak_DN_to_strong}, which requires more smoothness on $\Omega$ and boundary values.

We have
\begin{equation}
\Lambda_\gamma^s (v) = \gamma |\nabla u|^{p-2}\doo_\nu u,
\end{equation}
from which by expanding $\nabla u$ we get the equation
\begin{equation*}
\gamma |\nabla_T u + \nabla_\nu u|^{p-2}\doo_\nu u = \Lambda_\gamma^s (v).
\end{equation*}
Taking absolute values and reorganising we have
\begin{equation*}
(|\nabla_T u|^2 + |\nabla_\nu u|^2)^{\frac{p-2}{2}}|\doo_\nu u| = \frac{1}{\gamma}|\Lambda_\gamma^s (v)|.
\end{equation*}
We write the left hand side as a function of $|\doo_\nu u| = |\nabla_\nu u|= t$:
\begin{equation}
F(t) = (|\nabla_T u|^2 + t^2)^{\frac{p-2}{2}}t = \frac{1}{\gamma}|\Lambda_\gamma^s (v)|.
\end{equation}

There is a unique $t_0 \in [0,\infty[$ with $F(t_0) = \frac{1}{\gamma}|\Lambda_\gamma^s (v)|$.
To see this, observe that $\lim_{t \to 0} F(t) = 0$, $\lim_{t \to \infty} F(t) = \infty$, and that the function $F$ is strictly increasing and continuous.

Returning to the original formulation and expanding $|\nabla u|^2$ we therefore get
\begin{equation}
\gamma (|\nabla_T u|^2 + t_0^2)^\frac{p-2}{2} \doo_\nu u = \Lambda_\gamma^s (v),
\end{equation}
from which we can solve $\doo_\nu u$. We have proven the following lemma:
\begin{lemma}\label{lemma:doo_nu}
Suppose $\Omega$ is a bounded open set with $C^{1,\beta}$ boundary.
Let the boundary voltage $v \in C^{1,\beta}$ be known.
Then, for any boundary point $x_0 \in \doo \Omega$, we can compute $\doo_\nu u (x_0)$ from the strong DN map with the following algorithm:
\begin{enumerate}
\item Solve $t$ from the equation \[ (|\nabla_T u (x_0)|^2 + t^2)^{\frac{p-2}{2}}t = \frac{1}{\gamma(x_0)}|\Lambda_\gamma^s (v)(x_0)|. \]
\item Set $\doo_\nu u (x_0) = t\sign(\Lambda_\gamma^s (v)(x_0))$.
\end{enumerate}
\end{lemma}

The following remark is not used in this paper, but in general locating the points where $\nabla u = 0$ is interesting for $p$-harmonic functions $u$.
\begin{remark}
We know the boundary points $x_0$ where $\nabla u(x_0) = 0$, since $\doo_\nu u(x_0) = 0$ if and only if $\Lambda_\gamma^s (v)(x_0) = 0$, and the Dirichlet data determines $\nabla_T u$ at boundary points.
\end{remark}

We still need to recover the strong DN map from the weak one.
\begin{lemma}\label{lemma:weak_DN_to_strong}
Suppose that $\Omega$ has $C^{2,\beta}$ boundary, boundary values~$v \in C^{2,\beta}(\doo \Omega)$ and that $\nabla \gamma$ is H{\"o}lder-continous.
Then we can recover the pointwise values of the strong DN map
\begin{equation}
\gamma(x_0) \left|\nabla u(x_0)\right|^{p-2}\doo_\nu u (x_0)
\end{equation}
from the weak DN map (see definition~\ref{definition:DN_map})
\begin{equation}
\left\langle \Lambda_\gamma^w(v), g \right\rangle = \int_\Omega \gamma |\nabla u|^{p-2} \nabla u \cdot \nabla \tilde g \dx{x}.
\end{equation}
\end{lemma}
\begin{proof}
For $x_0 \in \doo \Omega$ define test functions~$\eta_\delta \colon \R^d \to \R$ by
\begin{equation}
\eta_\delta (x) = c_d \delta^{-d} (\delta - |x-x_0|)\chi_{B(x_0,\delta)}
\end{equation}
where $c_d > 0$ is selected so that $\int_{L} \eta_\delta \dx{x} = 1$,
where $L$ is the tangent space of $\doo \Omega$ at $x_0$.
Then $\eta_\delta$ lies in all first order Sobolev spaces with $|\nabla \eta_\delta(x)| = c_d \delta^{-d}$ in $B(x_0,\delta)$, and $0$ elsewhere.

Suppose $\nabla u(x_0)=0$.
By H{\"o}lder continuity of $|\nabla u|^{p-1}$ we get
\begin{align*}
\left|\langle \Lambda_\gamma^w (v),\eta_\delta\rangle\right| &\lesssim \| |\nabla u|^{p-2} \nabla u \cdot \nabla \eta_\delta \|_{L^1(\Omega \cap B(x_0,\delta))} \\
&\lesssim \| |\nabla u|^{p-1} |\nabla \eta_\delta| \|_{L^1(B(x_0,\delta))} \lesssim \|\delta^{\beta_1} \delta^{-d} \|_{L^1(\Omega \cap B(x_0,\delta))} \\
&\lesssim \delta^{\beta_1 - d} \delta^{d} = \delta^{\beta_1} \to 0 = \Lambda^s_\gamma(x_0)
\end{align*}
as $\delta \to 0$.

Suppose, then, that $\nabla u (x_0) \neq 0$.
Since $\nabla u_\eps \to \nabla u$ as $\eps \to 0$, there is $a > 0$ such that
\begin{equation}
|\nabla u_\eps (x_0)| \geq |\nabla u (x_0)|/2
\end{equation}
holds whenever $a > \eps \geq 0$.

We next prove that the equality
\begin{equation}\label{eq:a}
\gamma(x_0) \left(\left|\nabla u_\eps (x_0)\right|^2 + \eps \right)^\frac{p-2}{2}\doo_\nu u_\eps (x_0) = \lim_{\delta \to 0} \int_\Omega \gamma \gradeps^\frac{p-2}{2} \nabla u_\eps \cdot \nabla \eta_\delta \dx{x}
\end{equation}
holds.
By taking the limit $\eps \to 0$ we get the original claim if the rate of convergence in~\eqref{eq:a} is uniform in $\eps$.
By integrating by parts we get
\begin{align}
\smashoperator{\int_\Omega} \gamma &\gradeps^\frac{p-2}{2} \nabla u_\eps \cdot \nabla \eta_\delta \dx{x} \notag\\
=&-\smashoperator{\int_\Omega} \dive \left( \gamma \gradeps^\frac{p-2}{2}\nabla u_\eps \right) \eta_\delta \dx{x} \notag \\
&+ \smashoperator{\int_{\doo \Omega}} \eta_\delta \gamma \left(\left|\nabla u_\eps \right|^2 + \eps \right)^\frac{p-2}{2}\doo_\nu u_\eps \dx{S} \notag \\
= &\smashoperator{\int_{\doo \Omega \cap B(x_0,\delta)}} \eta_\delta \gamma \left(\left|\nabla u_\eps \right|^2 + \eps \right)^\frac{p-2}{2}\doo_\nu u_\eps \dx{S}, \label{int1}
\end{align}
since $u_\eps$ solves the perturbed equation,
and the integral~\eqref{int1} equals
\begin{align*}
\smashoperator{\int_{\doo \Omega \cap B(x_0,\delta)}} &\eta_\delta \gamma(x_0) \left(\left|\nabla u_\eps(x_0) \right|^2 + \eps \right)^\frac{p-2}{2}\nu(x_0) \cdot \nabla u_\eps(x_0) \dx{S} \\
+& \smashoperator{\int_{\doo \Omega \cap B(x_0,\delta)}} \eta_\delta \bigg( \gamma \left(\left|\nabla u_\eps \right|^2 + \eps \right)^\frac{p-2}{2}\doo_\nu u_\eps \\
&-\gamma(x_0) \left(\left|\nabla u_\eps(x_0) \right|^2 + \eps \right)^\frac{p-2}{2} \doo_\nu u_\eps(x_0) \bigg)\dx{S}\\
=&\left(\gamma(x_0) \left(\left|\nabla u_\eps (x_0)\right|^2 + \eps \right)^\frac{p-2}{2}\doo_\nu u_\eps (x_0) + o_{\delta \to 0}(1)\right)\smashoperator{\int_{\doo \Omega}} \eta_\delta \dx{S},
\end{align*}
since $\gamma$ and $\nabla u_\eps$ are continuous.
By normalisation of the test function this proves the equality~\eqref{eq:a}.

The rate of convergence is uniform in $\eps$ when $\delta$ is so small that $|\nabla u_\eps (x)|$ is bounded away from zero for $x \in B(x_0,\delta)$, since then the following $\eps$-indexed family of functions~\eqref{eq:equi} are equicontinuous with parameter $\eps$:
\begin{equation}\label{eq:equi}
x \mapsto \gamma(x) \left(|\nabla u_\eps(x)|^2 +\eps \right)^{\frac{p-2}{2}}\doo_\nu u_\eps (x).
\end{equation}
The family is H{\"o}lder-continuous on $\doo \Omega$ when $x - x_0$ is so small that $|\nabla u_\eps(x)| \geq m > 0$ for some $m$ and all $\eps < a$:
By lemma~\ref{lemma:eps_regularity} the functions $x \mapsto \nabla u_\eps (x)$ are equicontinuous with parameter $\eps$, and
the functions $w \mapsto \gamma \left( |w|^2 + \eps \right)^\frac{p-2}{2} w \cdot \nu$ are equicontinuous when $|w|$ is bounded from above and away from zero.
\end{proof}

\begin{remark}\label{remark:rellich}
We have shown that
\[ \int_\Omega (\doo_\alpha \gamma) |\nabla u|^p \dx{x} = \int_{\doo \Omega} \gamma (\alpha \cdot \nu) |\nabla u|^p \dx{S} - p\int_{\doo \Omega} (\doo_\alpha u) \Lambda_\gamma^s (u) \dx{S}. \]
Since we know $\gamma|_{\doo \Omega}$ by~\cite[theorem 1.1]{Salo:Zhong:2012}, $\nabla u|_{\doo \Omega}$ by lemma~\ref{lemma:doo_nu} and the strong DN map by lemma~\ref{lemma:weak_DN_to_strong}, we also know $\int_\Omega (\doo_\alpha \gamma) |\nabla u|^p \dx{x}.$
\end{remark}

\begin{remark}
In using the Rellich identity we need to know the values of the conductivity $\gamma$ on the entire boundary $\doo \Omega$.
The gradient $\nabla u_M$ should be large near $x_0$ and small elsewhere (we have not made this precise), so our proof is essentially local in nature.
\end{remark}

The main theorem~\ref{thm:main} now follows, since the value of the integral
\begin{equation}
\int_\Omega \doo_\alpha \gamma |\nabla u_M|^p \dx{x}
\end{equation}
is known (see remark~\ref{remark:rellich}) and it converges to $\doo_\alpha \gamma (x_0)$ by lemma~\ref{lemma:sz_main}.

\section*{Acknowledgements}
The research was partly supported by Academy of Finland.
Part of the research was done during a visit to the Institut Mittag-Leffler (Djursholm, Sweden).
The author is grateful to Mikko Salo and Joonas Ilmavirta for several discussions.

\bibliographystyle{plain}
\bibliography{math}

\end{document}